OpenStudy (anonymous):

write the equation of the line that has slope of 2/3 and passes through 1,3

OpenStudy (anonymous):

dont give me the answer. help me work through it please

OpenStudy (anonymous):

Ok do you know point slope form? y - y1 = m(x - 1)

OpenStudy (anonymous):

yes ive heard of it

OpenStudy (anonymous):

i think it starts off as y= 2/3x and then idk where to go from there

OpenStudy (anonymous):

Ok they give you a point and the slope so first you have to put in point slope form:) Fill in the values of x and y in the x1 and y1. and put the slope for m y - y1 = m(x - x1)

OpenStudy (anonymous):

y= 2/3x

OpenStudy (anonymous):

After you put it in point slope you can find slope-intercept form... y - 3 = 2/3(x - 1) Distribute

OpenStudy (anonymous):

now what do i do

OpenStudy (anonymous):

you distribute the 2/3 into the (x - 1) and what do get?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

ok.. to distribute you multiply the 2/3 by each factor in the parenthesis: 2/3 * x = ? 2/3 * -1 =?

OpenStudy (anonymous):

so its 0.666666666

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok so what is it

OpenStudy (anonymous):

2/3 * x = 2/3x

OpenStudy (anonymous):

sorry lost internet connection!

OpenStudy (anonymous):

can you give me the answer please